OpenStudy (anonymous):

solve for f 1/5=1/4+f

OpenStudy (zzr0ck3r):

subtract \(\frac{1}{4}\) from both sides

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

you are done

OpenStudy (anonymous):

oh

OpenStudy (zzr0ck3r):

what did you get for \(\frac{1}{5}-\frac{1}{4} \ = \ ?\)

OpenStudy (anonymous):

-1/20

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

f=-1/20

OpenStudy (anonymous):

thanks, wolfram makes things too confusing sometimes lol

OpenStudy (zzr0ck3r):

was it this \(\frac{1}{5}=\frac{1}{4}+f\) or \(\frac{1}{5}=\frac{1}{4+f}\)?

OpenStudy (anonymous):

the first one

OpenStudy (zzr0ck3r):

then you are good

OpenStudy (anonymous):

thanks :D